OpenStudy (anonymous):

If you were to place $2,500 in a savings account that pays 3% interest compounded continuously, how much money will you have after 5 years? Assume you make no other deposits or withdrawals. A. $2,904.59 B. $2,898.19 C. $431,078.73 D. $2,515.00

OpenStudy (anonymous):

@kc_kennylau @ranga do either of you know what type of formula I would use for this kind of question?

OpenStudy (ranga):

A = Pe^(rt)

OpenStudy (anonymous):

what do those terms stand for?

OpenStudy (ranga):

\[\Large A = Pe^{rt}\]

OpenStudy (anonymous):

I dont understand what numbers I would plug into which letters

OpenStudy (ranga):

A is the maturity amount P is the Principal r is the annual interest rate in decimals t is the time in years

OpenStudy (anonymous):

so it should be 2500=.03^5?

OpenStudy (ranga):

P = 2500 ; r = 0.03 ; t = 5 Find A.

OpenStudy (anonymous):

would it be 2500=x^.03(5) then?

OpenStudy (ranga):

A is NOT 2500 P is 2500 Find A

OpenStudy (anonymous):

x=2500^.03*5 ?

OpenStudy (ranga):

\[\Large A = 2500 \times e^{.03 * 5} = 2500 \times e^.15 = ?\]

OpenStudy (anonymous):

if that is correct then I got 3.23, if that formula isnt correct then I have no idea

OpenStudy (ranga):

you left out "e" in all your equations.

OpenStudy (anonymous):

NO! yes I just redid it & got 2904.59

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

thank you SO much, you're such a great help! good night :)

OpenStudy (ranga):

you are welcome ali.